# Deep Reinforcement Learning based Autonomous Decision-Making for Cooperative UAVs: A Search and Rescue Real World Application

### v1

@@ VI-D5 Discussion of Architectural Choices
VI-E Reward Function for the Guidance DRL
@@
VI-E1 Attractive Forces Towards the Target
@@
VI-E2 Repulsive Forces for Obstacle Avoidance
@@
  • When the obstacle is slightly to the left ($-\frac{\pi}{2}<\theta_{\text{obs}}<0$), a positive yaw adjustment is applied to turn the UAV right, away from the obstacle.
@@
  • The constants ±$\pm 0.5$ provide a base turning speed, and the term $\frac{\theta_{\text{obs}}}{\pi}$ adjusts the turning rate based on the angle, creating a smooth transition in yaw speed as the obstacle’s relative position changes.
@@ VI-E3 Calculating the reward
The reward is clipped to keep values between $-1.5$ and $1$, ensuring consistency with the bounds for task success and failure. Specifically, reaching a goal results in a reward of $+2$, while colliding with an obstacle incurs a penalty of $-2$. The final reward function is:
@@
VI-F Simulated Environment and Guidance Training
@@
VI-G GAT based Task Allocation Architecture
The TD3 DRL scheme will be used again for the task allocator, and the following architecture will be used as the actor-network. A similar architecture will be used for the critic networks with a concatenation that adds the actor’s action to the linear layers and again produces a single Q-value.
The task allocator for the UAV system is implemented using an GAT architecture known as GATConv [7], designed to determine the optimal node selection in a graph-based environment. The graph structure models the set of nodes representing locations/tasks to be visited by the UAVs, with each node being connected to all other nodes, forming a fully connected graph.
VI-G1 Graph Structure
The input graph is defined as follows:
  • Nodes: Each node represents a location or task that needs to be visited. The node features are encoded as a vector with three elements:
    • The first element indicates whether the node has already been visited (00 or $1$).
    • The second element indicates if the current UAV is at that node (00 or $1$).
    • The third element indicates if the other UAV is at that node (00 or $1$).
  • Edges: Each edge between nodes $i$ and $j$ carries a weight corresponding to the shortest unobstructed route between the two nodes. This route considers obstacles in the environment, providing a more accurate representation of the actual path that the UAV would need to take to travel between nodes $i$ and $j$.
VI-G2 GAT-Based Layer
The GATConv layer uses an attention mechanism to focus on the most relevant neighbouring nodes when making task allocation decisions. The attention mechanism assigns higher weights to edges representing shorter or more critical paths, based on the features of the nodes and the distance of the shortest unobstructed routes between them.
VI-G3 GAT-Based Model
The whole model is wrapped with a pre-processing layer and a post-processing layer. A fully connected layer proceeds the GATConv layers, expanding each node’s feature set to 64 with batch normalisation and a leakyReLU activation applied. There are two sequential GATConv layers, each with five attention heads. Following the GATConv layers, the data is fed to two fully connected layers to produce the final output.
The final output is a ten-dimensional vector representing the next node to visit. A softmax activation function is applied to this vector, converting the outputs into a probability distribution over the potential next nodes. The UAV selects the node with the highest probability as its next destination, guiding its movement through the environment. The layout of this model is shown in Fig. 6.
@@ VI-G3 GAT-Based Model
VI-G4 Decision Process
The GAT-based model processes the graph and generates an output that balances the need to visit unvisited nodes, the relative shortest paths between nodes, and the current locations of the cooperative UAVs (Two UAVs in our case). By incorporating the spatial and task-specific information directly into the graph, the network can dynamically allocate tasks to each UAV, optimising for efficiency and minimising redundant visits to already completed tasks.
VI-H Training of the GAT-Based Task Allocator
The training of the GAT-based task allocator is developed based on a DRL approach with a TD3 architecture. The training process aims to optimise the allocation of tasks among UAVs by learning to select the next node to visit based on the graph representation employed.
VI-H1 Training Setup
The training is performed on a workstation equipped with an NVIDIA RTX 3070m GPU [35], allowing for accelerated computation and efficient training. The model is trained over 20,000 episodes, where each episode consists of a complete attempt by the network to visit all nodes in the graph. During each episode, the UAVs interact with the environment by selecting nodes to visit, and the rewards guide the learning process toward optimal task allocation strategies.
VI-H2 Reward Function Design
The reward function is designed to encourage the UAVs to visit all nodes while minimising unnecessary travel and avoiding conflicts with other agents. It includes the following components:
  • Positive Rewards: The UAVs receive positive rewards for visiting previously unvisited nodes and for completing the task of visiting all nodes in the graph.
  • Negative Rewards: Penalties are applied for visiting nodes that have already been visited, nodes currently occupied by the other UAV, or nodes that are significantly distant from the current position. These penalties discourage inefficient task allocation and unnecessary travel.
  • Distance-Based Reward Adjustment: The total distance travelled by both UAVs to visit all nodes is also factored into the reward. A baseline distance is calculated by using a simple heuristic that selects the shortest edge to an unvisited node from each current node until the entire graph is visited. The actual distance travelled is measured against this baseline, and additional rewards or penalties are applied based on the relative efficiency of the UAVs’ paths.
This reward structure encourages the GAT-based task allocator to develop strategies that minimise travel distance while efficiently distributing the task of visiting nodes between the two UAVs. By rewarding the completion of the graph and penalising redundant or inefficient actions, the training process ensures that the network learns an effective policy for multi-agent task allocation.
VI-I Deploying the Software
VI-I1 AI-based Guidance AI
@@
The system utilised ROS 2 Isaac [37] to manage communication between the UAV’s sensors and control algorithms, facilitating seamless integration between hardware components and the guidance AI model. To interface the AI-generated guidance commands with the PX4 flight controller, the $\mu$XRCE-DDS client [39] was employed with the PX4_msgs ROS 2 package. This setup enabled efficient and reliable transmission of control commands from the AI model to the UAV, allowing for rapid and accurate execution of guidance decisions avoiding obstacles in real-world environments.
VI-I2 AI-based Task Allocation
@@
VII Test Results and Analysis
@@
VII-A Simulation Results
VII-A1 Validation of UAV AI-based Guidance
@@
VII-A2 Validation of GAT-based Task Allocation Strategy
@@
VII-B Testing in a Real-World Environment
@@
VII-B1 Real-World Sensor Output
@@
VII-B2 Obstacle Avoidance and Path Following
@@
[…]
@@
VII-B3 Real-World Task Allocation Performance
@@
VII-B4 LIDAR-SLAM Altitude Correction Performance
@@
VII-C The Sapience Autonomous Cooperative Drone Competition
@@
VII-C1 Review of System Performance in 3D Mapping Task
@@
VII-C2 Delivery Task System Performance
@@
VIII Additional Lessons Learned from the Deployment of the UAVs
@@
VIII-A Tuning the AI-based Guidance
@@
VIII-B Implementation of the Artificial Potential Field (APF)
@@
VIII-C Limitations of the Training Environment
@@
IX Conclusion
This work presents the development and deployment of a cooperative UAV system designed for SAR operations in GNSS-denied environments. The system integrates LIDAR-SLAM for real-time mapping and localisation, a DRL-based guidance AI for autonomous navigation, and a GAT-based task allocator for efficient multi-UAV coordination. These technologies are tested and validated within the NATO SAPIENCE Autonomous Cooperative Drone Competition, which provides a realistic and challenging platform for evaluating UAV performance in SAR-like scenarios.
The competition tasks—mapping, object detection, and synchronised deliveries—demonstrated the system’s capabilities in addressing critical challenges of SAR operations. The mapping task validates the ability to produce accurate 3D maps quickly, which is crucial for navigating hazardous environments. The delivery task showcases the system’s coordination and reliability in executing time-critical missions, such as synchronised aid delivery, emphasising the potential impact of cooperative UAV operations in life-saving applications.
Beyond SAR-specific applications, this work highlights a comprehensive cooperative UAV AI-based navigation, guidance, and task allocation system for GNSS-denied environments, validated through simulation and real-world testing. The integrated LIDAR-SLAM and a GAT-based task allocator demonstrate effective performance across diverse indoor settings, including complex structures with multiple obstacles, varying heights, and confined corridors. Real-world testing enables critical adjustments, such as implementing emergency obstacle avoidance mechanisms, refining height control with IMU-based velocity corrections, and addressing thermal management issues.
This system demonstrates a promising approach to autonomous UAV navigation and multi-agent task allocation, providing valuable insights into designing and deploying robust UAV systems for dynamic and collaborative operations in GNSS-denied indoor settings. Its performance in the SAPIENCE competition underscores the importance of advancing AI-based autonomous technologies for humanitarian applications. This work contributes to the broader goal of enabling faster, safer, and more efficient SAR operations and collaborative UAV missions in critical situations.

### v2

@@ VII-C6 Design Rationale
VII-D Reward Function for the Guidance DRL
@@
VII-D1 Attractive Forces Towards the Target
@@
VII-D2 Repulsive Forces for Obstacle Avoidance
@@
  • When the obstacle is slightly to the left ($\frac{\pi}{2}<\theta_{\text{obs}}<0$), a positive yaw adjustment is applied to turn the UAV right, away from the obstacle.
@@
  • The constants $\pm 0.5$ provide a base turning speed, and the term $\frac{\theta_{\text{obs}}}{\pi}$ adjusts the turning rate based on the angle, creating a smooth transition in yaw speed as the obstacle’s relative position changes.
@@ VII-D3 Calculating the reward
The reward is clipped to keep values between 1.5-1.5 and $1$, ensuring consistency with the bounds for task success and failure. Specifically, reaching a goal results in a reward of +2+2, while colliding with an obstacle incurs a penalty of 2-2. The final reward function is:
@@
VII-E Simulated Environment and Guidance Training
@@
VII-F GAT‑based Task Allocation Architecture
We cast cooperative task allocation as a single‑step, centralised RL problem trained with TD3. We adopted a central task allocator for two pragmatic reasons. (i) The Sapience Round‑1 rules explicitly specified a centralised decision maker, so aligning with that requirement avoided penalties or re‑engineering late in the cycle. (ii) Each UAV already maintains a single socket connection to the ground station for streaming sensor data for the mapping. Reusing that link to transmit task assignments was simpler and more reliable than implementing an additional peer‑to‑peer channel, reducing both software complexity and bandwidth overhead. The actor is a multi‑head Graph Attention Network (GATConv[8]); the two critics share the same graph encoder and concatenate the chosen action before the final linear layers.
VII-F1 Problem Formulation
At decision epoch $t$ the allocator observes a fully‑connected graph $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$. The vertex set $\mathcal{V}_{t}=\mathcal{V}^{\textsc{uav}}_{t}\cup\mathcal{V}^{\textsc{task}}_{t}$ contains UAV nodes ($|\mathcal{V}^{\textsc{uav}}_{t}|=N$) and task nodes ($|\mathcal{V}^{\textsc{task}}_{t}|=M$). Each node carries a 3‑D feature vector $\mathbf{x}_{v}=[v_{\text{visited}},\,v_{\text{self}},\,v_{\text{peer}}]^{\top}$ indicating whether the task is completed and which UAV (if any) occupies it. Edge features are the shortest collision‑free distances $d_{ij}$.
State
The global state is the graph embedding: $s_{t}=\langle\{\mathbf{x}_{v}\}_{v\in\mathcal{V}_{t}},\{d_{ij}\}_{(i,j)\in\mathcal{E}_{t}}\rangle$
Action
The allocator outputs an assignment matrix $a_{t}=\bigl[a_{nk}\bigr]\,\in\{0,1\}^{N\times(M+1)}$ where $a_{nk}=1$ assigns UAV $n$ to task $k$ (or to a dummy “idle” task $k=0$). Each UAV receives exactly one task: $\sum_{k}a_{nk}=1$.
Reward
Immediately after execution, we compute
$r_{t}=\underbrace{\Delta C_{t}}_{\text{tasks completed}}\;-\;\lambda_{d}\underbrace{\Delta D_{t}}_{\text{extra distance}}\;-\;\lambda_{c}\underbrace{C_{t}}_{\text{collision penalty}},$
where $\Delta C_{t}$ is the number of newly finished tasks, $\Delta D_{t}$ the incremental distance flown above a heuristic baseline, and $C_{t}$ a binary collision indicator. Positive reward is thus tied to productive, efficient, and safe allocations.
Terminal condition
Episodes terminate when all tasks are completed or when a fixed horizon $T$ is reached.
VII-F2 Network Details
A preprocessing linear layer expands node features to 64 dims (with BatchNorm and LeakyReLU). Two stacked GATConv layers, each with five attention heads, yield context‑aware node embeddings. A final MLP produces a (M+1)(M\!+\!1) -way softmax for every UAV, giving a joint action in a single forward pass (see Fig. 6). The same encoder is shared by the two critics, which regress separate Q‑values in the TD3 framework.
VII-F3 Training
We train for 20,000 episodes on an RTX 3070 Mobile GPU. Mini‑batches are drawn from a replay buffer; the critics follow clipped double‑Q learning while the actor is updated every other critic step, with target‑policy smoothing noise $\mathcal{N}(0,\sigma^{2})$ added to the next‑state actions.
VII-G Deploying the Software
VII-G1 AI-based Guidance AI
@@
[…]
VII-G2 AI-based Task Allocation
@@
VIII Test Results and Analysis
@@
VIII-A Simulation Results
VIII-A1 Validation of UAV AI-based Guidance
@@
Along with the APF-rewarded guidance AI, a model was also trained using a more standard distance-based reward, where the reward is given based on how much closer the drone moves to the goal per time-step. Several attempts were made with different hyperparameters and reward weightings, but the training could not produce a model suitable for testing that met the criteria used for the APF-rewarded AI guidance. Due to the time constraints imposed by the competition schedule and the one-week-long training run, the training could not be conducted until convergence.
@@
VIII-A2 Validation of GAT-based Task Allocation Strategy
@@
The results of this validation gave a model that always achieved success in four steps, meaning no unnecessary double moves. And achieved the lowest average distance travelled when compared to the median. This resulted in a score of -6.0m below the median distance travelled by the selected models.
To contextualise the proposed graph-attention task allocator, we compared its allocation outcome against representative classical baselines using the same waypoint-graph cost metric as our method (i.e., edge weights reflect corridor-constrained travel distance, and inter-node costs are shortest-path lengths on the graph). Specifically, we evaluated greedy nearest-task assignment, an auction-style bidding allocator (bids proportional to path cost), and an exact optimal one-step assignment (minimum-cost matching, recomputed after each completed task). On the field scenario used in this paper, all methods produced identical task sequences for both UAVs and therefore identical total travel distance and makespan. This indicates a ceiling effect arising from the small and highly structured competition task set: once true graph distances are used, the optimal allocation becomes largely unambiguous and can be recovered by several standard allocators. We therefore report these results as evidence that the proposed allocator is competitive on the deployed scenario, and we focus the remaining evaluation on deployment-relevant performance factors such as end-to-end autonomy, robustness, and real-world execution constraints.
TABLE I: Task allocation comparison of the GNN with classical methods in the Sapience layout
Method Total distance (m) Makespan (m) Steps
Greedy nearest-task 24.60 12.60 5
Auction / bidding 24.60 12.60 5
Optimal matching 24.60 12.60 5
Ours (GAT-DRL) 24.60 12.60 5
VIII-B Testing in a Real-World Environment
@@
VIII-B1 Real-World Sensor Output
@@
VIII-B2 Obstacle Avoidance and Path Following
@@
The testing environment was then expanded with the top half of the structure, including a corridor and a raised section. While the AI-based Guidance was able to change heights as needed, direct control of $z$-velocity initially caused oscillation due to overshooting. This was resolved by limiting $z$-velocity to $\pm 0.3\,\text{m/s}$, resulting in stable vertical movements. Additionally, a condition was added to ensure the UAV travelled fully over the raised ledge before descending, preventing rotor interference with the ledge.
@@
VIII-B3 Real-World Task Allocation Performance
@@
VIII-B4 LIDAR-SLAM Altitude Correction Performance
@@
VIII-C The Sapience Autonomous Cooperative Drone Competition
@@
VIII-C1 Review of System Performance in 3D Mapping Task
@@
VIII-C2 Delivery Task System Performance
@@
VIII-D Demonstration on an Alternative Layout
To provide additional evidence that the proposed cooperative framework is not specific to a single arena geometry, we include a brief demonstration on a secondary indoor layout used for system checks and demonstrations. This layout is intentionally simpler than the Sapience competition arena due to the complexities of erecting the Sapience arena and is therefore not presented as a stronger benchmark. Rather, it serves as a compact sanity check that the same end-to-end pipeline (waypoint graph construction, graph-based task allocation, and learned guidance execution) can be applied without altering the controller structure or task allocation logic.
Figure 23 shows the alternative layout and the set of task waypoints. The waypoint graph is constructed in the same manner as in the competition setting, and allocator decisions are computed using a GNN to determine the most efficient way to visit all nodes and complete the cooperative point cloud. The arena consists of two drones, one at each end, separated by a T-shaped obstacle in the middle with three mannequins arranged around it. The drones are also commanded to land at the opposite take-off site to reduce the time required to complete the task.
Table II compares the proposed GAT-DRL allocator against representative classical methods on the demonstration layout, using shortest-path distances on the same waypoint graph as the cost metric. The auction/bidding and optimal matching baselines yield identical performance in this instance, with a total travel distance of $26.30\,\mathrm{m}$ and a makespan of $14.88\,\mathrm{m}$ (where makespan denotes the maximum distance assigned to any single UAV). In contrast, the proposed allocator reduces the total travel distance to $21.19\,\mathrm{m}$ and the makespan to $11.94\,\mathrm{m}$, corresponding to reductions of $19.4\%$ and $19.8\%$, respectively. The proposed method also completes the allocation in fewer steps between nodes (4 versus 7), indicating a more stable and efficient division of work between the two UAVs on this layout.
Figure 22 summarises the resulting routes from the odometry as measured by the UAVs. Although this layout is simpler, the result demonstrates that the allocator and guidance policy remain coherent under a change in arena geometry, producing consistent task separations between agents and feasible collision-free routes.
Figure 22: Waypoint-graph visualisation of the Demo layout, showing the routes executed by each UAV.
Figure 23: The routes of the two UAVs when completing the Demo task.
To complement the top-down layout depiction, Fig. 24 shows the aggregated 3D point cloud collected during the Demo-layout flight.
Figure 24: The combined 3D point cloud collected from both UAVs as they have completed their allotted tasks.
TABLE II: Task allocation comparison of the GNN with classical methods in the Demo layout
Method Total distance (m) Makespan (m) Steps
Greedy nearest-task 29.26 14.64 7
Auction/bidding 26.30 14.88 7
Optimal matching 26.30 14.88 7
Ours (GAT-DRL) 21.19 11.94 4
IX Additional Lessons Learned from the Deployment of the UAVs
@@
IX-A Tuning the AI-based Guidance
@@
IX-B Implementation of the Artificial Potential Field (APF)
@@
IX-C Limitations of the Training Environment
@@
X Future Work
The three-stage NATO Sapience programme provides a natural roadmap for extending our framework beyond the present indoor demonstration.
X-A Outdoor and Mixed Indoor–Outdoor SAR
The second competition will be held entirely outdoors, followed by a combined indoor/outdoor finale. Operating without Global Navigation Satellite Systems (GNSS) in feature–sparse exteriors demands: (i) fully on–board autonomy—communication with the base station will be severed, so all high‐level decisions must arise from inter‐UAV links; (ii) richer odometry, e.g. visual–inertial–LiDAR fusion to maintain drift below $0.5\,\%$ of distance travelled; and (iii) ruggedised airframes capable of lifting heavier battery and sensor payloads ( 7kg MTOW).
X-B Multi–Agent Deep RL Guidance
The present work trains independent TD3 agents that share state vectors at inference time. Future research will explore Multi‐Agent DRL (MADRL) objectives that explicitly model pairwise collision avoidance and cooperative coverage through joint value functions or centralised critics.
X-C Decentralised Task Allocation
With the base station removed, our Graph Attention Network must execute on board. Two deployment topologies will be studied: a master–slave scheme in which a single UAV performs inference and broadcasts assignments, and a fully peer‐to‐peer consensus that amortises computation across the fleet and eliminates a single point of failure.
X-D Collaborative Perception and Mapping
We will extend the depth–IMU fusion that eliminates indoor LiDAR Z‐drift to a multi‐agent odometry pipeline. Sharing pose graphs and loop closures over a low‐bandwidth mesh could reduce accumulated vertical error by an order of magnitude, enabling altitude‐critical tasks such as window ingress.
These directions will culminate in a robust, decentralised multi-UAV SAR stack capable of seamless indoor–outdoor transition and long-range operation without external infrastructure.
Moving past the competition setting, several research directions will mature our framework into a general‑purpose indoor autonomy stack. First, we will replace the current waypoint-based navigation with goal-conditioned multi-agent DRL policies that learn cooperative, obstacle-aware trajectories directly from graph-structured sensor observations. This shift will allow each UAV to reason jointly about its neighbours and the environment in a unified latent space, eliminating the need for predefined intermediate goals.
Safety at close quarters will be bolstered through the real‑time exchange of lightweight perceptual cues—such as truncated LiDAR point clouds and depth key‑frames—so that each agent can augment its local occupancy map with the perspectives of its teammates. Shared perception is expected to reduce occlusion‑induced collisions and enable tighter formations in narrow corridors.
Accurate state estimation remains a bottleneck for large‑scale indoor operations. Building on our depth–IMU fusion, we will extend the filter to a fleet-level optimisation that incorporates inter-agent loop closures and ultra-wide-band ranging, yielding a globally consistent altitude and eliminating residual LiDAR-SLAM Z-drift.
Collectively, these avenues will transform the current prototype into a robust, scalable platform for collaborative UAV autonomy in cluttered, GNSS-denied environments.
XI Conclusion
This work presents an extensive real-world demonstration of a cooperative, DRL-guided UAV system operating in GNSS-denied indoor environments. We integrated LiDAR-SLAM for on-board localisation and mapping, a velocity-level DRL guidance policy shaped by an APF reward, and a centralised GAT-based allocator for multi-UAV coordination—then validated the complete stack under competition pressure, rather than in simulation alone.
The NATO SAPIENCE Autonomous Cooperative Drone Competition provided a realistic proxy for SAR deployments. Across mapping, object detection, and synchronised delivery tasks, our system showed it can: (i) rapidly produce actionable 3‑D maps of complex structures, (ii) coordinate time‑critical, multi‑agent behaviours, and (iii) sustain reliable autonomy amidst clutter, occlusions, and tight passages. These trials exposed—and allowed us to solve—practical hurdles that seldom appear in simulation studies, including emergency obstacle-avoidance fallbacks, thermal management, and altitude drift suppression via depth–IMU fusion.
By moving beyond constrained 2D controllers, small arenas, or purely waypoint-level actions, we demonstrate that DRL policies can scale to full-scale airframes, richer state-action spaces, and prolonged indoor missions. The resulting framework provides a concrete blueprint for deploying autonomous, cooperative UAV teams in life-saving SAR scenarios and other GNSS-denied operations.
Overall, our findings underscore a central message: closing the sim-to-real gap requires end-to-end field experimentation. The lessons, architectures, and reward designs reported here provide a tested foundation on which future work—both within and beyond SAPIENCE—can build toward faster, safer, and more effective collaborative aerial robotics.